A motor boat traveled 12 miles with the current, turned around, and returned 12 miles against the current to its starting point. The trip with the current took 2 hours and the trip against took 3 hours. Find the speed of the boat and the speed of the current.
Found 2 solutions by stanbon, deborabr:
Answer by stanbon(75887)

A motor boat traveled 12 miles with the current, turned around, and returned 12 miles against the current to its starting point. The trip with the current took 2 hours and the trip against took 3 hours. Find the speed of the boat and the speed of the current.
---------------------------
With current DATA:
distance = 12 miles ; time = 2 hrs ; rate = d/t = 12/2 = 6 mph
----
Against current DATA:
distance = 12 miles ; time = 3 hrs ; rate = d/t = 12/3 = 4 mph
------
Equation:
b + c = 6 mph
b - c = 4 mph
-------
2b = 10
b = 5 mph (speed of the boat in still water)
----
Solve for "c":
b + c = 6
5 + c = 6
c = 1 mph (speed of the current)
=======================================
Cheers,
Stan H.

Answer by deborabr(181)

speed of the boat
12/2
6 miles per hour
speed of the current
the boat against the current 12/3 = 4 miles per hour
6 - 4 = 2 miles per hour is the speed of the current